### v3

# Rethinking Machine Unlearning for Large Language Models

Abstract
We explore machine unlearning (MU) in the domain of large language models (LLMs), referred to as LLM unlearning. This initiative aims to eliminate undesirable data influence (e.g., sensitive or illegal information) and the associated model capabilities, while maintaining the integrity of essential knowledge generation and not affecting causally unrelated information. We envision LLM unlearning becoming a pivotal element in the life-cycle management of LLMs, potentially standing as an essential foundation for developing generative AI that is not only safe, secure, and trustworthy, but also resource-efficient without the need of full retraining. We navigate the unlearning landscape in LLMs from conceptual formulation, methodologies, metrics, and applications. In particular, we highlight the often-overlooked aspects of existing LLM unlearning research, e.g., unlearning scope, data-model interaction, and multifaceted efficacy assessment. We also draw connections between LLM unlearning and related areas such as model editing, influence functions, model explanation, adversarial training, and reinforcement learning. Furthermore, we outline an effective assessment framework for LLM unlearning and explore its applications in copyright and privacy safeguards and sociotechnical harm reduction.
1 Introduction
Large language models (LLMs) have shown exceptional proficiency in generating text that closely resembles human-authored content. However, their ability to memorize extensive corpora may also lead to ethical and security concerns. These include societal biases and stereotyping (Bender et al., 2021; Motoki et al., 2023; Kotek et al., 2023), the generation of sensitive, private, harmful, or illegal content (Nasr et al., 2023; Wen et al., 2023; Karamolegkou et al., 2023; Patil et al., 2024), ease of jailbreaking (Wei et al., 2023; Zou et al., 2023; Liu et al., 2023a), and possible malicious use in developing cyberattacks or bioweapons (Barrett et al., 2023; Hendrycks et al., 2023). These concerns emphasize the need to adeptly and efficiently tailor pre-trained LLMs to suit diverse safety contexts while meeting specific requirements of users and sectors.
With the costly and prolonged training periods of LLMs, retraining these models to eliminate undesirable data effects is often impractical. Machine unlearning (MU) has emerged as an alternative to remove the influence of undesirable data and associated model capabilities from the pre-trained models (Cao & Yang, 2015; Bourtoule et al., 2021; Nguyen et al., 2022; Si et al., 2023; Zhang et al., 2023a; Eldan & Russinovich, 2023; Yao et al., 2023). For example, MU is used as a strategy to prevent the generation of copyrighted material from the Harry Potter series in (Eldan & Russinovich, 2023). In the context of classification tasks, MU has been extensively studied by Ginart et al. (2019); Neel et al. (2021); Ullah et al. (2021); Sekhari et al. (2021); Golatkar et al. (2020); Jia et al. (2023). Yet, its application and understanding in LLMs remains limited, where models are typically used for generative tasks such as summarization, sentence completion, paraphrasing, and question answering. Therefore, this paper specifically concentrates on exploring the MU problems in LLMs, termed ‘LLM Unlearning’.
As data-model scales continue to grow, the emergence of LLM unlearning introduces new challenges and complexities, as will be elaborated on in Sec. 2. For example, current research efforts (Lu et al., 2022; Jang et al., 2022; Ilharco et al., 2022; Eldan & Russinovich, 2023; Wu et al., 2023b; Yu et al., 2023; Zhang et al., 2023c; Yao et al., 2023) suffer from a lack of standardized corpora and evaluation methods for LLM unlearning, often facing issues related to unlearning precision and efficiency. We also refer readers to Sec. 3 for a summary of existing LLM unlearning tasks.
Figure 1: Demonstration of how MU can be incorporated into LLM development cycle. The landscape of LLM unlearning will be mainly navigated from applications (‘why’), methods (‘where’ and ‘how’), and evaluations.
Although preliminary surveys of LLM unlearning have been provided by Si et al. (2023) and Zhang et al. (2023a), this paper is, to the best of our knowledge, the first to offer a thorough and in-depth analysis of LLM unlearning. Our contributions are summarized below. Fig. 1 provides an overview of the LLM unlearning landscape that we explore.
(1) Surveying: We conduct an in-depth review of the foundational concepts and principles of LLM unlearning, delving into the problem formulation, categories of unlearning methods, evaluation approaches, and practical applications.
(2) Uncovering: We bring to light previously overlooked dimensions of LLM unlearning, e.g., emphasizing the significance of precisely defining the unlearning scope, elucidating the interplay between data and model interactions, and exploring the adversarial assessment of unlearning efficacy.
(3) Connecting: We establish connections between LLM unlearning and other relevant problems and domains, providing a comparative analysis with related topics such as model editing, influence function, and adversarial learning.
(4) Forecasting: We offer insights into the future of LLM unlearning by identifying novel opportunities & challenges.
This paper is positioned to reassess the challenge of LLM unlearning, refining its scope across various dimensions: conceptual formulation (Sec. 3), methods (Sec. 4), assessment (Sec. 5), and applications (Sec. 6); see the schematic overview in Fig. 1. We delve into each dimension through surveying, uncovering, connecting, and forecasting. We conclude that unlearning will be a valuable tool for making LLMs more trustworthy, but making more progress on this will require updating the unlearning paradigm. We aspire for this work to pave the way for developing LLM unlearning, illuminating its opportunities, challenges, and untapped potential.
2 Preliminaries and Related Work
LLM unlearning has garnered attention for addressing trustworthiness concerns such as toxicity (Lu et al., 2022), copyright and privacy (Jang et al., 2022; Eldan & Russinovich, 2023; Wu et al., 2023b), fairness (Yu et al., 2023), model editing (Ilharco et al., 2022; Zhang et al., 2023c), hallucination (Yao et al., 2023), and sensitive knowledge (Barrett et al., 2023; Hendrycks et al., 2023). In this section, we present a succinct overview of MU, tracing its journey from traditional ML models to the emerging challenges in LLMs.
MU for non-LLMs.
The study of MU can be traced back to non-LLMs in response to data protection regulations such as ‘the right to be forgotten’ (Cao & Yang, 2015; Hoofnagle et al., 2019; Bourtoule et al., 2021; Nguyen et al., 2022). Due to its capability of assessing data influence on model performance, the landscape of MU has expanded to encompass diverse domains, such as image classification (Ginart et al., 2019; Golatkar et al., 2020; Neel et al., 2021; Ullah et al., 2021; Sekhari et al., 2021), text-to-image generation (Gandikota et al., 2023; Zhang et al., 2023b; Kumari et al., 2023; Fan et al., 2024), federated learning (Liu et al., 2020; Wang et al., 2022; Che et al., 2023; Liu et al., 2023b; Halimi et al., 2022), and graph neural networks (Chen et al., 2022; Chien et al., 2022; Wu et al., 2023a).
In the literature, ‘exact’ unlearning, which involves retraining the model from scratch after removing specific training data points, is often considered the gold standard. However, this approach comes with significant computational demands and requires access to the entire training set (Thudi et al., 2022). To address these challenges, many research efforts have shifted towards the development of scalable and effective approximate unlearning methods (Golatkar et al., 2020; Warnecke et al., 2021; Becker & Liebig, 2022; Thudi et al., 2022; Jia et al., 2023; Chen et al., 2023a). In addition, probabilistic methods with certain provable removal guarantees have been explored, often leveraging the concept of differential privacy (Ginart et al., 2019; Guo et al., 2019; Neel et al., 2021; Ullah et al., 2021; Sekhari et al., 2021).
Challenges of MU for LLMs.
LLM unlearning introduces new challenges and complexities.
First, LLMs are trained on massive amounts of data, which can unintentionally introduce biases and the memorization of personal and confidential information. Accordingly, it becomes challenging to precisely define and localize the ‘unlearning targets’, such as the subset of the training set or a knowledge concept that needs to be removed. Therefore, current studies on LLM unlearning (Lu et al., 2022; Jang et al., 2022; Ilharco et al., 2022; Eldan & Russinovich, 2023; Wu et al., 2023b; Yu et al., 2023; Zhang et al., 2023c; Yao et al., 2023) are typically context and task-dependent. There is a lack of standardized corpora for LLM unlearning.
Second, the growing size of LLMs and the rise of black-box access to LLM-as-a-service present challenges in developing scalable and adaptable MU techniques to LLMs (Bucknall & Trager, 2023; Casper et al., 2024a). This also affects performance evaluation, given the absence of retraining as a benchmark. To address these challenges, previous studies have proposed approaches like in-context unlearning (Pawelczyk et al., 2023) and fictitious unlearning (Maini et al., 2024), where the former enables unlearning on black-box models, and the latter provides alternatives to retraining.
Third, the scope of unlearning is often underspecified for LLMs. This issue is similar to challenges faced in model editing (Mitchell et al., 2022). For instance, effective unlearning should ensure that LLMs delete knowledge of the targeted data within the predefined scope while simultaneously maintaining its utility for data outside of this scope.
Fourth, despite the potential of LLM unlearning in diverse applications, there is a notable absence of comprehensive and reliable evaluation. For example, recent studies (Shi et al., 2023; Patil et al., 2024) have demonstrated that sensitive information can be reverse-engineered from an edited model, even if efforts were made to delete this information. This highlights the need for thorough and adversarial evaluations and the design of more mechanistic methods to guarantee the authenticity of unlearning.
3 Unpacking LLM Unlearning
In this section, we formalize the LLM unlearning problem and delve into its configuration. We will outline the inherent facts and complexities in its formulation.
Problem statement.
In light of the literature on unlearning (Bourtoule et al., 2021; Jia et al., 2023; Kurmanji et al., 2023), and its progression in LLMs (Pawelczyk et al., 2023; Yao et al., 2023; Ishibashi & Shimodaira, 2023; Maini et al., 2024), we define the problem of LLM unlearning below.
(LLM unlearning) How can we efficiently and effectively eliminate the influence of specific ‘unlearning targets’ and remove associated model capabilities while preserving model performance for non-targets?
We dissect the above statement from the perspectives: (1) unlearning targets, (2) influence erasure, (3) unlearning effectiveness, and (4) efficiency. See Table 1 for a summary of existing LLM unlearning studies based on (1)-(4).
Table 1: A summary of existing LLM unlearning problems through unlearning targets, influence erasure, effectiveness, and efficiency. An asterisk () indicates the incapability of evaluating unlearning for LLMs due to the impracticality of retraining these models.
Related work
Unlearning targets/tasks
Influence erasure methods
Effectiveness:
(I) In-scope evaluation for unlearning efficacy
(O) Out-of-scope evaluation for model utility
Efficiency
(Lu et al., 2022)
Reducing toxic content,
avoiding undesirable sentiments,
and preventing repeated text generation
Reward-reinforced model fine-tuning
(I) Toxic prompts, specific sentiments,
& repetitive sentences
(O) Unlearning target-irrelevant prompts
N/A
(Jang et al., 2022)
Degenerating private information,
w/ unlearning response irrelevant to this info
Gradient ascent-based fine-tuning
(I) Prompts from training data extraction
(O) Natural language understanding tasks
Runtime cost
(Kumar et al., 2022)
Text de-classification, w/ unlearning
response close to that of retraining
Sharded, isolated, sliced, and
aggregated (SISA) training via adapter
(I) No evaluation for unlearning efficacy
(O) Test set
Runtime cost
Memory cost
(Ilharco et al., 2022)
(Zhang et al., 2023c)
Degenerating toxic content
Task vector-based parameter-efficient
fine-tuning via LoRA
(I) Prompts leading to toxic generation
(O) Perplexity on other datasets
N/A
(Wang et al., 2023)
Text de-classification/de-generation,
unlearning specific words in translation,
w/ response close to that of retraining
KL-divergence-based fine-tuning
(I) Training subset
(O) Test set
Runtime cost
(Yu et al., 2023)
Unlearning gender and profession bias,
with de-biased unlearning response
Weight importance-informed
& relabeling-based fine-tuning
(I) Biased prompts
(O) No evaluation for model utility
N/A
(Pawelczyk et al., 2023)
Text de-classification, w/ unlearning
response close to that of retraining
In-context learning
(I) Training subset
(O) Retain & test sets
Black-box access
(Eldan & Russinovich, 2023)
Degenerating Harry Potter-related
book content, w/ unlearning response
irrelevant to Harry Potter
Relabeling-based fine-tuning
(I) Questions and their rephrased/hard versions
about Harry Potter
(O) NLU tasks
N/A
(Ishibashi & Shimodaira, 2023)
Unlearning knowledge from QA dataset,
with refusal response (e.g., ‘I don’t know’)
Relabeling-based fine-tuning
(I) Adversarial and original questions
about forgotten knowledge
(O) Other QA prompts
N/A
(Chen & Yang, 2023)
Text de-classification and de-generation,
with response close to that of retraining
KL divergence-based parameter-
efficient fine-tuning via adapter
(I) Training subset
(O) Retain & test sets
Runtime cost
(Wu et al., 2023b)
Degenerating private information,
w/ unlearning response irrelevant to this info
Importance-based neuron editing
(I) Memorized private data points
(O) Test set
Runtime cost
(Yao et al., 2023)
Degenerating harmful prompts,
degenerating Harry Potter-related
book content, and reducing hallucination
Integration of gradient ascent,
random labeling,
& KL divergence-based fine-tuning
(I) Prompts related to unlearning targets
(O) NLU tasks
Runtime cost
(Maini et al., 2024)
Unlearning biographical knowledge
about fictitious authors.
Fine-tuning with various objectives
(I) Q&A about the unlearning authors
(O) Q&A about other authors and general facts
Runtime cost
(Patil et al., 2024)
Degenerating sensitive information
using factual information as a testbed
Model editing techniques and
constrained finetuning
(I) Prompts for unlearned factual knowledge
(O) Prompts for unrelated factual knowledge
White-box v.
black-box access
$\bullet$ Unlearning targets: Unlearning tasks may take on various forms and are closely related to the unlearning objectives. For instance, one could focus on data influence removal, while the other could emphasize model capability removal. Although these two aspects are intertwined, the former is often crucial for intellectual property (IP) protection, while the latter is more practical for AI alignment and safety. The literature identifies unlearning targets as specific data points, which could involve content containing harmful, unethical, or illegal language (Jang et al., 2022; Wu et al., 2023b). They have also been represented by higher-level unlearned knowledge, expressed through an unwanted text prompt or concept (Lu et al., 2022; Yao et al., 2023; Eldan & Russinovich, 2023). For example, Eldan & Russinovich (2023) defined the unlearning target as ‘Harry Potter’-related content, with the objective to avoid generating such content irrespective of where the content was learned: from the copyrighted material, blog posts, or news articles.
$\bullet$ Influence erasure: Erasing the influence of unlearning targets and associated model capabilities requires a joint examination of both data and model influences rather than studies in isolation. Specifically, it is important to scrutinize the contributions of data sources to undesired model outputs, as well as the roles played by individual components within a model in generating these undesirable outcomes. This dual examination allows us to gain a more comprehensive understanding of the mechanisms driving these outputs, thereby facilitating the development of unlearning strategies to prevent them effectively. The objective of achieving complete influence erasure also implies the importance of robustness and generalization in unlearned behavior. When evaluating LLM unlearning, especially when using approximate methods shown in Table 1, a rigorous criterion is needed. Patil et al. (2024) underscore this viewpoint by demonstrating that sensitive information can persist in model weights after editing or unlearning, and by Maini et al. (2024) and Shi et al. (2023) who expose the limitations of conventional unlearning methods.
$\bullet$ Unlearning effectiveness: The effectiveness of LLM unlearning extends beyond merely diminishing the specific data influence or model capability. A crucial aspect of effectiveness is the concept of unlearning scope, as inspired by the editing scope (Mitchell et al., 2022). This has implications for the development of metrics to evaluate the effectiveness of unlearning and the design of effective unlearning methods. The unlearning scope defines the success of influence erasure by its capacity to accurately modify the model’s behavior for in-scope examples, which are either directly or indirectly related to unlearning targets, such as paraphrased text prompts to be degenerated (Zhong et al., 2023; Patil et al., 2024). At the same time, it must also ensure the generation consistency for out-of-scope examples, necessitating the preservation of model generation capabilities post-unlearning. Differentiating between in-scope and out-of-scope examples for unlearning is its own unsolved problem, as it requires determining when facts logically imply one another (Hase et al., 2023b; Cohen et al., 2023).
$\bullet$ Unlearning efficiency & feasibility: Current research efforts focused on the development of rapid unlearning methods for pre-trained LLMs due to the significant re-training cost (Jang et al., 2022; Eldan & Russinovich, 2023; Yao et al., 2023). However, beyond computational efficiency, LLMs present additional efficiency challenges. These include the complexity and, at times, the infeasibility of pinpointing and attributing training data points designated for unlearning. Additionally, there is the challenge of executing unlearning in the context of black-box LLMs (Pawelczyk et al., 2023), where interactions with models are constrained to input-output queries.
Distinguishing LLM unlearning from conventional MU.
According to the above dimensions, LLM unlearning involves a broader range of targets, which are often context-dependent and less clearly defined. Moreover, the effectiveness of LLM unlearning is not limited to forgetting the influence of specific data points but also includes defining a broader unlearning scope for model capability removal. Furthermore, there is a critical need to devise more mechanistic methods that guarantee effective and robust unlearning, while also enhancing their practicality and feasibility.
Mathematical modeling.
Building upon the high-level formulation of LLM unlearning problems presented earlier, we next provide mathematical modeling details and discuss the associated design choices. To facilitate comprehension, we provide a commonly-used formulation of LLM unlearning problems below. While this may not be the sole or optimal approach to LLM unlearning, it incorporates several key elements that are essential to the problem setup.
$\begin{array}[]{l}\displaystyle\min_{\bm{\theta}}\,\,\underbrace{\mathbb{E}_{(x,y_{\mathrm{f}})\in\mathcal{D}_{\mathrm{f}}}[\ell(y_{\mathrm{f}}|x;\bm{\theta})]}_{\text{Forget}}+\lambda\underbrace{\mathbb{E}_{(x,y)\in\mathcal{D}_{\mathrm{r}}}[\ell(y|x;\bm{\theta})]}_{\text{Retain}}\end{array}$ (2)
where $\ell(y|x;\bm{\theta})$ denotes the prediction loss of using $\bm{\theta}$ given the input $x$ with respect to the response $y$, $\mathcal{D}_{\mathrm{f}}$ and $\mathcal{D}_{\mathrm{r}}$ refer to ‘forget’ and ‘retain’ sets which will be explained later, $y_{\mathrm{f}}$ denotes the desired model response post-unlearning, and $\lambda\geq 0$ is a regularization parameter to balance ‘forget’ and ‘retain’ (e.g., $\lambda=0$ if retain set is not given a priori).
In the dataset setup of LLM unlearning, we typically assume access to a forget set ($\mathcal{D}_{\mathrm{f}}$), the influence of which should be eliminated in LLM generation. For instance, $\mathcal{D}_{\mathrm{f}}$ might consist of a collection of harmful or toxic prompt-response pairs designated for degeneration (Yao et al., 2023). Moreover, if the original training set is available, then $\mathcal{D}_{\mathrm{f}}$ can be composed of a subset of training data points related to the unlearning target. Alternatively, it can be generated using synthesized data points based on a higher-level unlearned knowledge concept, or it can be derived from a set of extracted training data points reverse-engineered from the given LLM itself. In practice, the forget set $\mathcal{D}_{\mathrm{f}}$ is not required to belong precisely to the LLM’s training corpus. Given the diversity and extensive size of the model’s training data, the content we aim to unlearn is more likely to represent a general concept. Thus, LLM unlearning needs to not only unlearn specific training samples but also generalize to similar samples that share common characteristics.
Besides the forget set $\mathcal{D}_{\mathrm{f}}$, sometimes there is a need for a retain set ($\mathcal{D}_{\mathrm{r}}$), which contains samples that are not subject to unlearning and used to preserve the utility of the unlearned model. Through the lens of the unlearning scope we discussed earlier, the forget set ($\mathcal{D}_{\mathrm{f}}$) provides in-scope examples earmarked for unlearning, while the retain set ($\mathcal{D}_{\mathrm{r}}$) involves examples out of the unlearning scope. Some recent studies have also attempted to develop LLM unlearning approaches that operate independently of access to forget and/or retain sets (Pawelczyk et al., 2023; Li et al., 2023b).
We next introduce the model and optimization setups for LLM unlearning. Unlearning is performed at the post-model training phase. As shown in (2), a common unlearning objective is to efficiently update the original pre-trained model so that the updated model can unlearn on $\mathcal{D}_{\mathrm{f}}$ while retaining its generation capability on $\mathcal{D}_{\mathrm{r}}$. In addition, another design element is unlearning response ($y_{\mathrm{f}}$), referred to as the response of an unlearned model to in-scope examples. For example, in the stateful LLM unlearning method aimed at erasing information related to ‘Who’s Harry Potter?’ (Eldan & Russinovich, 2023), the unlearning response is based on word replacements using generic translations, like substituting ‘Quidditch’ with ‘Skyball’, as part of the unlearning process. However, this type of approach may inadvertently blur the line between LLM hallucination and legitimate responses, highlighting the need for improvements in unlearning response design. Another choice is empty response (Wu et al., 2023b; Patil et al., 2024), given by the reject ‘I don’t know’ (Patil et al., 2024) or the customized response by ‘masking’ the unlearning information using specialized tokens such as $\blacksquare$’ in (Wu et al., 2023b). However, we need to ensure that the empty response targets only examples within the unlearning scope. Otherwise, frequent rejections may occur, potentially diminishing the user experience with LLMs.
4 LLM Unlearning Methods Revisited: Prior Art and Overlooked Principles
In this section, we begin with a review of existing MU (machine unlearning) methodologies adapted for LLM unlearning. Following that, we delve into the design factors that are often overlooked in the literature, including the intertwined data-model influence, the relationship with model editing, and adversarial training.
Review of existing unlearning principles.
Existing LLM unlearning methods can be broadly categorized into two groups: model-based and input-based. Model-based methods involve modifying the weights and/or architecture components of LLMs to achieve the unlearning objective (Jang et al., 2022; Lu et al., 2022; Yao et al., 2023; Yu et al., 2023; Chen & Yang, 2023; Zhang et al., 2023c; Hase et al., 2023a; Wu et al., 2023b; Rafailov et al., 2023), e.g., following the mathematical formulation in Sec. 3. Input-based methods design input instructions (Madaan et al., 2022; Zheng et al., 2023; Pawelczyk et al., 2023), such as in-context examples or prompts, to guide the original LLM (without parameter updating) towards the unlearning objective.
In the literature, the predominant research emphasis lies on model-based methods as shown in Tab. 1. We review some of the most representative approaches below.
$\bullet$ Gradient ascent and its variants (Jang et al., 2022; Yao et al., 2023; Yu et al., 2023): Gradient ascent (GA) stands as one of the most straightforward unlearning methods, updating the model parameters by maximizing the likelihood of mis-prediction for the samples within the forget set $\mathcal{D}_{\mathrm{f}}$ (Jang et al., 2022). However, it is worth noting that GA alone can be sensitive to the choice of hyperparameters during optimization, such as the number of ascent steps and the learning rate (Jia et al., 2023; Fan et al., 2024). This has given rise to improved variants of GA. For example, Yao et al. (2023) incorporate random labeling to augment the unlearning objective and ensure utility preservation on the retain set $\mathcal{D}_{\mathrm{r}}$.
Another variant of GA is to transform it into a gradient descent approach, which aims to minimize the likelihood of predictions on relabeled forgetting data (Yu et al., 2023). This gradient descent-based fine-tuning, over relabeled or randomly labeled forgetting data, is also employed in (Eldan & Russinovich, 2023), where generic translations are used to replace the unlearned texts. GA and its variants often involve fine-tuning pre-trained LLMs for unlearning purposes. To enhance efficiency, parameter-efficient fine-tuning (PEFT) techniques could be employed. Chen & Yang (2023) fine-tune an adapter over the unlearning objective that acts as an unlearning layer within the LLM. Zhang et al. (2023c) use LoRA to create task vectors and accomplish unlearning by negating tasks under these task vectors.
$\bullet$ Localization-informed unlearning (Meng et al., 2022; Yu et al., 2023; Wu et al., 2023b): The pursuit of parameter efficiency is also in line with the objective of identifying and localizing a subset of model units (e.g., layers, weights, or neurons) that are essential for the unlearning task. For example, the process of localization can be accomplished through representation denoising, also known as causal tracing, in (Meng et al., 2022), focusing on the unit of model layers. Previous work shows that it is important to delete information about unlearning targets wherever it is represented in models in order to protect against attacks (Patil et al., 2024). In addition, gradient-based saliency (Yu et al., 2023) is employed to identify the crucial weights that need to be fine-tuned to achieve the unlearning objective. In (Wu et al., 2023b), neurons that respond to unlearning targets are identified within the feed-forward network and subsequently selected for knowledge unlearning. In the context of vision models, unlearning can also benefit from localizing weights salient to unlearning, as demonstrated by Jia et al. (2023) and Fan et al. (2024). Furthermore, the concept of localization-informed unlearning resonates with the future modular machine learning solution development (Menik & Ramaswamy, 2023). This modularity allows the emerging foundation models to be partitioned into manageable subparts, facilitating easier maintenance and independent updates for each component.
$\bullet$ Influence function-based methods: While the influence function (Koh & Liang, 2017; Bae et al., 2022) is a standard approach to assess the effect of data removal on model performance (Izzo et al., 2021; Warnecke et al., 2021), it is not commonly employed in the context of LLM unlearning for two main reasons: the computational complexity involved in inverting the Hessian matrix, and the reduced accuracy resulting from the use of approximations in influence function derivation (Jia et al., 2023). However, the potential of influence functions in LLM unlearning may be underestimated, given that scalability issue has been improved by Grosse et al. (2023), and approximation errors can be mitigated by focusing on localized weights that are salient to unlearning, as described in the previous category.
$\bullet$ Other model-based methods: Other types of LLM unlearning methods fall outside the categories mentioned above. For example, Jang et al. (2022); Chen & Yang (2023) show that sequential unlearning outperforms batch unlearning However, as indicated by Gu et al. (2024), sequential editing of LLMs may compromise their general capabilities. Therefore, the study of sequential unlearning in LLMs continues to be an unresolved issue.
$\bullet$ Input-based vs. model-based: Input-based strategies (Madaan et al., 2022; Zheng et al., 2023; Pawelczyk et al., 2023) have also shown promise in addressing the challenges posed by the restricted access to black-box LLMs and achieving parameter efficiency of LLM unlearning. Here the learnable parameters are given by input prompts rather than model weights/architecture components. However, we posit that input-based methods may not necessarily yield genuinely unlearned models, leading to weaker unlearning strategies compared to model-based methods because modifying the inputs of LLMs alone may not be sufficient to completely erase the influence of unlearning targets (Toyer et al., 2023). This assertion is also supported by the existence of hard or even adversarial in-scope examples associated with unlearning targets and the challenge to remove their influence in LLMs (Zhong et al., 2023; Patil et al., 2024).
Exploring data-model interactions.
A key objective of unlearning is to eliminate the influence of the forgotten data points/knowledge on the model’s performance. However, this process is not studied in isolation: It is closely connected to exploring the influence of model weights or architecture components. Unlearning requires a sense of locality, which involves addressing the specific unlearning target and its associated unlearning scope. Consequently, exploring model influence helps identify the specific, localized areas of the model that are relevant to this locality. This is further reinforced by the surveyed weight localization techniques (Meng et al., 2022; Yu et al., 2023; Wu et al., 2023b; Patil et al., 2024). Thus, model influence and data influence are intertwined in LLM unlearning, and a comprehensive understanding of the former can streamline the process of handling data influence.
Relationship with model editing.
Model editing, closely related to LLM unlearning, focuses on the local alteration of pre-trained models’ behavior to introduce new knowledge or rectify undesirable behaviors. First, the objective of editing could align with that of unlearning when editing is introduced to erase information. Second, like unlearning scope, editing scope (Mitchell et al., 2022; Hase et al., 2023b; Cohen et al., 2023) is crucial to ensure that editing is executed without compromising the generative capabilities of the model outside the defined scope. Third, both model editing and unlearning can be approached using the ‘locate first, then edit/unlearn’ principle. Localization in the context of model editing has also been applied to various elements, including neurons (Dai et al., 2021), network layers (Meng et al., 2022; Gupta et al., 2023), and feed-forward components of LLMs (Geva et al., 2020; Li et al., 2023c).
Despite the aforementioned connections, there are clear distinctions between LLM unlearning and editing. First, the unlearning response is sometimes unknown compared to the editing response. The specificity of an incorrect or improper unlearning response could be seen as a form of LLM hallucination after unlearning. Second, although unlearning and model editing may share some common algorithmic foundations, the former does not create new answer mappings. Rather, its central aim is the comprehensive elimination of the influence attributed to a specific knowledge or concept within a pre-trained LLM. Third, we can differentiate model editing from unlearning from the perspective of ‘working memory’. It is known in (Li et al., 2022) that working memory in LLMs is maintained through neuron activations rather than weight-based long-term memory. Thus, existing memory-based model editing techniques (Li et al., 2022; Mitchell et al., 2022; Madaan et al., 2022; Zheng et al., 2023) focus on updating short-term working memory instead of altering the long-term memory encapsulated in the model’s weights. Yet, we posit that unlearning requires more mechanistic approaches that facilitate ‘deep’ modifications to pre-trained LLMs.
Adversarial training for robust unlearning.
An increasing body of research highlights the weaknesses of existing unlearning methods (Shi et al., 2023; Maini et al., 2024), particularly in their vulnerability to test-time adversaries attempting to jailbreak unlearned models. This issue has been explored by Patil et al. (2024) for LLMs and Zhang et al. (2023d) for diffusion models, and inspires us to integrate adversarial training (Madry et al., 2017) into the unlearning process, resulting in what we term adversarial unlearning. However, this approach has received relatively little attention thus far. While adversarial unlearning increases training costs, it also presents new opportunities. For instance, localization-informed unlearning can significantly reduce the computation expenses associated with adversarial unlearning by focusing on a small portion of model units for updating. Furthermore, advanced adversarial training techniques such as fast adversarial training (Shafahi et al., 2019; Wong et al., 2020; Zhang et al., 2022) and generalized adversarial training (Zhu et al., 2019; Kumari et al., 2019; Zhang et al., 2022; Robey et al., 2023; Casper et al., 2024b) have the potential to enhance the scalability of adversarial unlearning while preserving its effectiveness.
Reinforcement learning.
The mainstream technique for aligning LLMs with human values is RLHF (reinforcement learning from human feedback) and its variants (Christiano et al., 2017; Ouyang et al., 2022; Bai et al., 2022; Yuan et al., 2023; Lee et al., 2023; Rafailov et al., 2023; Casper et al., 2023). However, RLHF is resource-intense: (1) it requires human inputs that are expensive to collect, and (2) it is computationally costly (i.e., the standard three-stage aligning procedure). LLM unlearning arises as an alternative aligning method, where collecting negative (i.e., low-quality and harmful) samples is much easier through user reporting or (internal) red teaming than positive (i.e., high-quality and helpful) samples which often require hiring humans. Furthermore, we can envision a reinforced unlearning paradigm with a properly defined reward function for the unlearned tasks. In a similar vein, Rafailov et al. (2023); Lu et al. (2022) utilize a reward model to assist in unlearning.
5 Assessment
In this section, we discuss key considerations for designing an effective evaluation pipeline for LLM unlearning.
Tasks and benchmarks.
Datasets related to harmful content degeneration, personal identification information removal, and copyrighted information prevention serve as suitable benchmarks for evaluating the effectiveness of LLM unlearning. Some notable examples of these datasets include: The Enron dataset, which comprises employee emails publicly disclosed during Enron’s legal investigation by the Federal Energy Regulatory Commission (Wu et al., 2023b), the Training Data Extraction Challenge dataset used in (Jang et al., 2022), the Harry Potter book series dataset (Eldan & Russinovich, 2023; Shi et al., 2023), the toxicity generation dataset (Lu et al., 2022; Gehman et al., 2020), and the TOFU dataset for fictitious unlearning (Maini et al., 2024).
Evaluation of unlearning effectiveness.
We examine efficacy from three perspectives: comparison with retraining (i.e., the gold standard of unlearning), ‘hard’ in-scope evaluation or robustness, and training data detection.
$\bullet$ LLM unlearning vs. retraining. In classic unlearning paradigms (Golatkar et al., 2020; Thudi et al., 2022; Jia et al., 2023; Fan et al., 2024), retraining a model from scratch after removing the forgotten data from the original training set is regarded as exact unlearning. However, the scalability challenges of retraining LLMs make it difficult to establish a performance upper bound for evaluating LLM unlearning. A recent solution by Maini et al. (2024) is to incorporate fictitious data (synthetic author profiles) into the model training paradigm. Since the injected set never appeared in the original pretraining set, LLM fine-tuning can simulate the retraining process over the newly-introduced set. Despite the progress with regard to the specialized TOFU dataset (Maini et al., 2024), there is still a pressing need to advance current evaluation metrics and pipelines to accurately assess the gap between (approximate) LLM unlearning methods and exact unlearning.
$\bullet$ ‘Hard’ in-scope evaluation or robustness. As demonstrated in Sec. 3, unlearning is generally context and task-dependent, contingent upon an unlearning scope. Another effectiveness metric of LLM unlearning is to ensure forgetting concerning in-scope unlearned examples, even for those ‘hard’ ones that fall within the unlearning scope but may not be directly associated with the unlearning targets. The assessment of ‘hard’ in-scope examples can be achieved by techniques such as paraphrasing what LLMs intend to forget or creating multi-hop questions (Zhong et al., 2023). Evaluating ‘hard’ in-scope examples aligns seamlessly with the underlying principles of ‘worst-case’ or ‘adversarial’ evaluation methods for unlearning, resembling LLM jailbreaking attacks (Zhang et al., 2023d; Yong et al., 2023; Patil et al., 2024). For instance, it is shown by Yong et al. (2023) that unlearning a scope using an English-only example would not guarantee a similar unlearned outcome when translated into other languages. It is also crucial to evaluate the robustness of unlearned LLMs after fine-tuning. Recent studies have revealed that fine-tuning LLMs can sometimes lead to the re-emergence of behaviors that were not anticipated (Yang et al., 2023; Qi et al., 2023; Lermen et al., 2023; Yong et al., 2023).
$\bullet$ Training data detection or membership inference. Membership inference attacks (MIA) (Shokri et al., 2017), designed to detect if a data point is part of a victim model’s training set, serve as a crucial data privacy-unveiled metric for evaluating machine unlearning methods (Thudi et al., 2022; Jia et al., 2023). This metric gains even more significance in the context of LLM unlearning, particularly when retraining is not an option. This concept is also connected to training data memorization (Carlini et al., 2022), as well as training data extraction attacks (Nasr et al., 2023) in LLMs. In the realm of LLM unlearning, privacy-related evaluation metrics have been explored and considered in various studies (Shi et al., 2023; Wu et al., 2023b; Jang et al., 2022; Pawelczyk et al., 2023; Maini et al., 2024).
Utility preservation.
Another crucial metric is to ensure the retained generation capabilities of unlearned LLMs on standard language modeling tasks that fall outside the unlearning scope. For example, evaluation on natural language understanding tasks (Jang et al., 2022; Eldan & Russinovich, 2023; Yao et al., 2023) and perplexity (Ilharco et al., 2022; Zhang et al., 2023c) has been considered in the literature. In line with evaluating the effectiveness of LLM unlearning on ‘hard’ in-scope examples, it is equally crucial to assess utility preservation using ‘hard’ out-of-scope examples, achieved by e.g., using data transformations. Lastly, we note that it can be difficult to determine the exact scope for some unlearning target (Hase et al., 2023b; Cohen et al., 2023), so part of the challenge here is deciding which generation capabilities should be retained in the first place.
Efficiency.
Computation cost has been a predominant efficiency metric when evaluating LLM unlearning methods, as demonstrated in Table 1. In addition to that, efforts have been made to extend LLM unlearning to black-box models, without access to model parameters, as demonstrated by Pawelczyk et al. (2023). Furthermore, memory efficiency could also serve as a crucial efficiency metric. The distinction from parameter efficiency is that current parameter-efficient fine-tuning methods still impose substantial memory costs for storing LLMs and for executing back-propagation (Malladi et al., 2023). Thus, a future research direction is to explore memory-efficient fine-tuning methods for LLM unlearning.
6 Applications of LLM Unlearning
In this section, we discuss two application categories facilitated by LLM unlearning: the first focused on data influence and the second on model capabilities.
Copyright and privacy protection.
One application of unlearning involves legal and ethical considerations around the fair use of training data. Algorithmic disgorgement is the term applied in law and policy for the requirement put on a company by a regulator, such as the Federal Trade Commission (FTC) in the United States, to completely destroy a model that was trained on data without legal consent (Li, 2022; Goland, 2023; Belkadi & Jasserand, 2023; Achille et al., 2023). The most famous case to-date is the FTC calling for the destruction of a weight loss application by WW International, whose underlying model contained illegal health information from children. Unlearning presents a viable alternative to complete disgorgement by removing the effect of the illegal data.
Also, the tension between data owners (e.g., authors) and LLM service providers is escalating, leading to legislation such as legal disputes involving OpenAI, Meta, and New York Times (Small, 2023; Grynbaum & Mac, 2023). This trend is likely to persist due to increasing societal concerns about AI data usage. The need for copyright-protected content removal aligns with the capabilities of LLM unlearning. However, it is often challenging to pinpoint the exact sources of training data that need to be deleted, giving rise to the issue of data attribution (Li et al., 2023a). For example, the leakage related to the ‘Harry Potter’ series (Eldan & Russinovich, 2023) can have multiple possible causes, e.g., the books were used in the LLM’s training data, the training data containing online discussions related to the series, or the LLM using retrieval-augmented generation (Gao et al., 2023) which might lead to leakage from the search results.
Similar to deleting copyrighted information from the training data, another scenario is preventing LLMs from leaking user privacy, especially personal identification information (PII). This concern is closely related to LLM memorization and training data extraction (Carlini et al., 2019, 2021, 2022; Jang et al., 2022; Nasr et al., 2023).
Sociotechnical harm reduction.
Another application of LLM unlearning is alignment (Ouyang et al., 2022), aimed at aligning LLMs with human instructions and making sure generated text conforms to human values. Unlearning can be used to forget harmful behaviors such as the production of toxic, discriminatory, illegal, or morally undesirable outputs (Shevlane et al., 2023), e.g., instructions to build CBRN (chemical, biological, radiological, and nuclear) weapons. Unlearning, as a safety alignment tool, can happen at the different stages of LLM development, e.g., before, during, or after alignment. Current research has focused on the ‘pre-alignment’ stage (Yao et al., 2023), there may be untapped opportunities in the others.
Hallucinations, which involve the generation of false or inaccurate content that may appear plausible, are a significant challenge in LLMs. Previous research has demonstrated that unlearning can reduce LLM hallucinations by targeting and unlearning factually incorrect responses given specific questions (Yao et al., 2023). Since hallucination is likely to be caused by multiple sources, the possible usage is to unlearn factually incorrect data that serve as the source of commonly shared hallucinations or misconceptions.
LLMs are known to generate biased decisions and outputs (Perez et al., 2022; Tamkin et al., 2023; Cui et al., 2023). In the vision domain, unlearning has proven to be an effective tool for reducing discrimination to enable fair decision-making (Sattigeri et al., 2022; Chen et al., 2023b). In the language domain, unlearning has been applied to mitigate gender-profession bias (Yu et al., 2023) and many other fairness issues (Sattigeri et al., 2022; Oesterling et al., 2023; Kadhe et al., 2023). However, more opportunities exist, such as unlearning stereotypes in training data.
LLMs are also known to be vulnerable to jailbreaking attacks (Wei et al., 2023; Qi et al., 2023; Huang et al., 2023) (i.e., adversarially crafted prompts that lead the LLM to generate undesired outputs) as well as poisoning/backdoor attacks (Rando & Tramèr, 2023; Carlini et al., 2023). Unlearning can be a natural solution for both types of attacks given the existing success of unlearning as a defense against adversarial attacks in other domains (Wang et al., 2019; Li et al., 2021; Liu et al., 2022; Jia et al., 2023).
7 Discussion and Conclusion
This position paper rethinks the paradigm of unlearning for modern LLMs to uncover its under-explored aspects. To achieve this, we dissect LLM unlearning into four essential aspects: formulation, methodologies, evaluation metrics, and applications. We show that there are considerable challenges in both foundational research and practical, use case-driven research. These include: (Generality) A desired solution for LLM unlearning should take into account the generality of the unlearning target and dataset choice, accommodate various model setups including both white-box and black-box scenarios, and consider the specifics of the unlearning method. (Authenticity) LLM unlearning should focus on effectively removing both data influence and specific model capabilities, in order to authenticate unlearning across a range of evaluation methods, particularly in adversarial contexts. (Precision) LLM unlearning should precisely define the scope of unlearning, while ensuring the preservation of general language modeling performance outside this unlearning scope.
By examining the current state of the art, we gain insights for the future development of LLM unlearning. For example, localization-informed unlearning shows promise with possible dual advantages of efficiency and efficacy. Effective unlearning requires careful consideration of data-model influences and adversaries. Despite similarities between LLM unlearning and model editing, they differ in their formulation and methodological design. Furthermore, insights gained from the study of LLM unlearning could catalyze technological advancements in other types of foundation models, e.g., large vision-language models.
8 Broader Impacts
The broader impact of our work, particularly its ethical implications and societal consequences, revolves around the responsible use of data and maintaining the integrity of large language models (LLMs) and other foundational AI models. Our research aims to explore machine unlearning methods to enhance safety and security, uphold user privacy, reduce biases, and guarantee the generation of trustworthy information by AI systems. The methodologies and frameworks we explored have the potential to significantly influence ethical AI practices in the future. We recognize, however, that the societal implications of progress in machine unlearning are intricate and diverse. This necessitates ongoing dialogue and scrutiny, especially in relation to the dynamic field of AI ethics and governance. As machine unlearning finds application in practical scenarios, it is essential to thoughtfully address any ethical dilemmas and societal ramifications that may arise. This underscores our dedication to advancing AI technology in a manner that is both responsible and mindful of its broader impacts.
9 Acknowledgements
S. Liu, J. Jia, and Yuguang Yao were supported by the National Science Foundation (NSF) Robust Intelligence (RI) Core Program Award IIS-2207052 and the Cisco Faculty Research Award. P. Hase and M. Bansal were supported by NSF-CAREER Award 1846185, NSF-AI Engage Institute DRL-2112635, DARPA MCS Grant N66001-19-2-4031, and Google PhD fellowship. We also extend our gratitude to the MIT-IBM Watson AI Lab for their support in this project.